# Completeness of closed Kleinian flat Pseudo-Riemannian Manifolds of Signature (2,2)

### v1

@@ 1.2.1. Reduction result
Let $G\leq{\operatorname{SL}}(d,\mathbb{R})$ be a semisimple Lie subgroup, and fix a Cartan decomposition $G=KA^{+}K$, where $K\leq G$ is a maximal compact subgroup and $A^{+}$ is the exponential of a closed positive Weyl chamber (see Section 3.1). For a subgroup $G^{\prime}\leq G$, denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ with respect to this decomposition. Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$. We say that $A^{+}(G^{\prime})$ does not contract a subspace $F\subset\mathbb{R}^{d}$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $g\in A^{+}(G^{\prime})$ and all $x\in F$.
Fix $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces. We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$ (in particular $p\geq\tfrac{d}{2}$). We prove the following result.
@@
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a discrete group acting freely, properly discontinuously, and cocompactly on $\Omega$. Suppose that the linear part $L(\Gamma)$ of $\Gamma$ is contained in a semisimple Lie group $G\leq{\operatorname{SL}}_{d}(\mathbb{R})$ such that:
@@
    there exists $\frac{d}{2}\leq p\leq d$ such that $G$ preserves an orbit $\mathcal{F}\subset\operatorname{Gr}_{p}(\mathbb{R}^{d})$ whose closure consists of pairwise transverse subspaces;
@@
    there exists $H\in\mathcal{F}$ such that $A^{+}(L(\Gamma))$ does not contract $H$.
Then there exists $H^{\prime}$ in the closure of $\mathcal{F}$ that is $L(\Gamma)$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $H^{\prime}$.
This proposition generalizes a result of Tholozan [Tho15, Proposition 2.4], where the statement is proved in the case $G=\mathrm{U}(n,1)$ and $\mathcal{F}$ is the orbit of a complex hyperplane. In our case, the set of isotropic planes in $\mathbb{R}^{2,2}$ has two connected components, each of which is an orbit under the identity component ${\operatorname{SO}}_{0}(2,2)$; moreover, the elements of each orbit are pairwise transverse (see Proposition 3.7). Let us briefly explain this well known fact. Write $\mathbb{R}^{2,2}=\mathbb{R}^{2}\oplus\mathbb{R}^{2}$, endowed with the nondegenerate quadratic form of signature $(2,2)$ given by $g\oplus(-g)$, where $g$ is the standard Euclidean metric. The graph of any linear isometry $A\in O(2)$ is a totally isotropic plane in $\mathbb{R}^{2,2}$. One can show that the set of isotropic planes in $\mathbb{R}^{2,2}$ is exactly $O(2)$; in particular, it has two connected components. The key observation is that if $A,B\in{\operatorname{SO}}(2)$ are distinct elements, then their graphs intersect only at $0\in\mathbb{R}^{2,2}$ and so they are transverse. Indeed, their intersection coincides with the set of fixed vectors of $AB^{-1}$, and a nontrivial element of ${\operatorname{SO}}(2)$ has no nonzero fixed vectors. This property no longer holds in ${\operatorname{SO}}(n)$ for $n>2$, and it is precisely at this point that the signature $(2,2)$ plays a crucial role. Nevertheless, we believe that Proposition B is of independent interest and may apply in other settings, which deserve further investigation. For instance, one may consider the case where $G={\operatorname{SO}}(4,\mathbb{C})$, as well as the case where $G$ is a rank-one irreducible simple Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$. These cases are part of an ongoing project of the three authors.
@@ 1.2.2. Quotient geometry
Let $G$ be a homothety Lie group, and let $G_{\theta}=R\ltimes_{\theta}G$ be a semidirect product, where $R$ is a linear Lie group and $\theta$ commutes with a nontrivial homothety of $G$. Let $\pi\colon G_{\theta}\to R$ be the natural projection. Let $\Gamma\leq G_{\theta}$ be a discrete subgroup, and let $H=\overline{\pi(\Gamma)}$. Set $\Gamma_{nd}:=\Gamma\cap\pi^{-1}(H^{\circ})$. Then $\Gamma_{nd}$ admits a nilpotent syndetic hull $S$ in $G_{\theta}$ with the property that $\overline{\pi(\Gamma_{nd})}=\overline{\pi(N)}=H^{\circ}$.
@@ 2.2. Algebraic and cohomological constraints
Let $\Omega\subset\mathbb{R}^{d}$ be an open set divided by a discrete subgroup $\Gamma\leq{\operatorname{SL}}(d,\mathbb{R})\ltimes\mathbb{R}^{d}$. Then $\operatorname{cd}(\Gamma)\geq d$.
@@ 3. Reduction of the flat affine geometry
3.1. Cartan projection
An important feature of the proof of Proposition 3.1 is the Cartan projection. Let $G$ be a semisimple Lie group, and denote by $\mathfrak{g}$ its Lie algebra. Let $K\leq G$ be a maximal compact subgroup and $\mathfrak{h}$ be the orthogonal of the Lie algebra $\mathfrak{k}$ of $K$ in $\mathfrak{g}$} for the Killing form. A Cartan subspace $\mathfrak{a}$ of $\mathfrak{g}$} is a maximal abelian subalgebra of $\mathfrak{h}$.
The restricted Weyl group $W$ of $G$ is the quotient of the normalizer of $\mathfrak{a}$ in $K$ (for the adjoint action) by the centralizer of $\mathfrak{a}$ in $K$. A fundamental system of $\mathfrak{a}$ for the action of $W$ is called a closed Weyl chamber. It is a closed convex cone with nonempty interior. Given such a closed Weyl chamber $\overline{\mathfrak{a}}^{+}$, we denote by $A^{+}$ the exponential of $\overline{\mathfrak{a}}^{+}$ ; the group $G$ then admits the following Cartan decomposition
$G=KA^{+}K.$
Given an element $g\in G$, we can thus write $g=kak^{\prime}$ according to this decomposition. The elements $k,k^{\prime}\in K$ are not unique, but the element $a\in A^{+}$ is. The Cartan projection of $g$ is the unique element $X\in\overline{\mathfrak{a}}^{+}$ such that $a=\exp(X)$.
@@ 3.2. General reduction result
For a semisimple Lie subgroup $G\leq{\operatorname{SL}}(d,\mathbb{R})$, we write $G=KA^{+}K$ for a chosen Cartan decomposition of $G$ (in the notation of Section 3.1 above). For a subgroup $G^{\prime}\leq G$, we denote by $A^{+}(G^{\prime})$ the set of Cartan projections of elements of $G^{\prime}$ in this decomposition.
Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ that is invariant under the maximal compact subgroup $K$. We say that a subset $X$ of $G$ does not contract $F$ with respect to $|\cdot|$ if $F$ is invariant under all elements of $X$ and if $|g\cdot x|\geq|x|$ for all $g\in X$ and all $x\in F$. In what follows, we will simply say that $X$ does not contract $F$. To prove Proposition 3.1, we first establish the stronger result given by Proposition B, inspired by arguments from [Tho15]. To fix notation, let $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces, note that the linear action of $\mathrm{SL}(d,\mathbb{R})$ induces a natural action on $\operatorname{Gr}_{p}(\mathbb{R}^{d})$. We say that $E,F\in\operatorname{Gr}_{p}(\mathbb{R}^{d})$ are transverse if $E+F=\mathbb{R}^{d}$.
The proof of Proposition B requires some preparation. Throughout this section, we work under the assumptions of Proposition B. Write $\Omega=\Gamma\cdot\mathcal{C}$ with $\mathcal{C}\subset\Omega$ compact. Let $\varepsilon>0$ be such that $\mathcal{C}+B_{\varepsilon}\subset\Omega$, where $B_{\varepsilon}$ denotes the Euclidean ball of radius $\varepsilon$ centered at 0, defined with respect to the Euclidean norm $\lvert\cdot\rvert$.
@@
Let $y\in\partial\Omega$. Then there exist sequences $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$, $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$, and $(H_{n})_{n\in\mathbb{N}}\subset\mathcal{F}$ such that, up to extraction of a subsequence:
@@
    for every $n$, one has $(H_{n}+g_{n}\cdot x_{n})\cap B(g_{n}\cdot x,\varepsilon)\subset\Omega$;
@@
    $H_{n}$ converges in $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ to some element $H_{y}$ in the closure of $\mathcal{F}$, with $(H_{y}+y)\cap B(y,\varepsilon)\subset\partial\Omega$, where $B(y,\varepsilon)$ is Euclidean ball of radius $\varepsilon$ centered at $y$.
@@
Since $\Omega=\Gamma\cdot\mathcal{C}$, there exist sequences $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$ and $(g_{n})_{n\in\mathbb{N}}\subset\Gamma$ such that $g_{n}\cdot x_{n}\to y$. This proves the first point. For all $n\in\mathbb{N}$, we have
@@
For each $n$, write the Cartan decomposition $L(g_{n})=w_{n}a_{n}w_{n}^{\prime}$, with $w_{n},w_{n}^{\prime}\in K$ and $a_{n}\in A^{+}(L(\Gamma))$. Let $H$ be the $p$-subspace appearing in the hypothesis of Proposition B, and set $H_{n}:=w_{n}\cdot H$. We claim that $L(g_{n})\cdot B_{\varepsilon}$ contains $H_{n}\cap B_{\varepsilon}$. Indeed, let $v\in w_{n}\cdot H\cap B_{\varepsilon}$ and set $x=w_{n}^{-1}v\in H$. Since $K$ preserves the Euclidean norm $\lvert\cdot\rvert$ and $a_{n}$ does not contract $H$, we have $\lvert a_{n}^{-1}x\rvert\leq\lvert x\rvert=\lvert v\rvert\leq\varepsilon$. Therefore $u:=w_{n}^{\prime-1}(a_{n}^{-1}x)$ belongs to $B_{\varepsilon}$, and $L(g_{n})u=w_{n}a_{n}w_{n}^{\prime}\bigl(w_{n}^{\prime-1}a_{n}^{-1}x\bigr)=w_{n}x=v$. Hence $w_{n}\cdot H\cap B_{\varepsilon}\subset L(g_{n})\cdot B_{\varepsilon}$, which proves the claim.
It follows from (1) that $\Omega$ contains $y_{n}+(H_{n}\cap B_{\varepsilon})$, establishing the second point. By compactness of $\overline{\mathcal{F}}$, we may assume (up to extracting a subsequence) that $H_{n}$ converges in the Grassmannian to some $H_{y}\in\overline{\mathcal{F}}$. Then $\overline{\Omega}$ contains $y+(H_{y}\cap B_{\varepsilon})$, and since $y\in\partial\Omega$, the inclusion must in fact be
$(H_{y}\cap B_{\varepsilon})+y\subset\partial\Omega.$
@@
In the previous lemma, we do not use the transversality of the elements of the closure of $\mathcal{F}$. In particular, the lemma asserts that each point of the boundary contains a small neighborhood of an affine subspace.
The next step is to show the uniqueness of the subspace $H_{y}$. Before doing so, we record the following elementary fact about transversality.
@@
Let $y\in\partial\Omega$, and let $H_{y}$ be an element of the closure of $\mathcal{F}$ as in Lemma 3.2. Then $H_{y}$ is the unique element of $\overline{\mathcal{F}}$ with the property that there exists a neighborhood $\mathcal{V}$ of $y$ in $\Omega$ such that $(y+H_{y})\cap\mathcal{V}\subset\partial\Omega$.
@@
Let $(y_{n})_{n\in\mathbb{N}}$, $(H_{n})_{n\in\mathbb{N}}$, and $H_{y}$ be as in Lemma 3.2. By shrinking $\varepsilon$ if necessary, we may assume that $B(y,\varepsilon)\subset\mathcal{V}$.
We proceed by contradiction and assume the existence of a subspace $H\in\overline{\mathcal{F}}$, distinct from $H_{y}$, that satisfies the hypothesis of the lemma. Then $H_{y}$ and $H$ are transverse. Since $y_{n}\to y$ and $H_{n}\to H_{y}$, and using Lemma 3.4, we deduce that, for $n$ large enough, the set $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)$ intersects $(H+y)\cap B(y,\varepsilon)$.
However, by Lemma 3.2, we have $(H_{n}+y_{n})\cap B(y_{n},\varepsilon)\subset\Omega$, while $(H+y)\cap B(y,\varepsilon)\subset(H+y)\cap\mathcal{V}\subset\partial\Omega$, a contradiction. ∎
@@
Let $y\in\partial\Omega$, and let $H_{y}$ be as in Proposition 3.2. Then $H_{y}+y\subset\partial\Omega$.
@@
Let $\mathcal{U}$ be the set of points $z\in H_{y}+y$ such that $(H_{y}+y)\cap B(z,\varepsilon)\subset\partial\Omega$. We will show that $\mathcal{U}$ is both open and closed (and nonempty) in $H_{y}+y$, so that $\mathcal{U}=H_{y}+y$, which will complete the proof.
Clearly, the set $\mathcal{U}$ is closed and contains $y$. We now show that $\mathcal{U}$ is also open. Let $z\in\mathcal{U}$ and $z^{\prime}\in(H_{y}+y)\cap B(z,\varepsilon/2)$. Then $(H_{y}+y)\cap B(z^{\prime},\varepsilon/2)\subset\partial\Omega$. Thus, by Lemma 3.5, we have $H_{y}=H_{z^{\prime}}$. By the definition of $H_{z^{\prime}}$, this implies $H_{y}\cap B(z^{\prime},\varepsilon)=H_{z^{\prime}}\cap B(z^{\prime},\varepsilon)\subset\partial\Omega,$ hence $z^{\prime}\in\mathcal{U}$. Therefore $\mathcal{U}$ is open, which completes the proof. ∎
@@
First, we claim that two affine subspaces contained in $\partial\Omega$ whose linear parts lie in $\overline{\mathcal{F}}$ must be parallel. By Lemma 3.6, there exists at least one affine subspace contained in $\partial\Omega$ whose linear part belongs to $\overline{\mathcal{F}}$.
@@
Assume, for contradiction, that there exist two affine subspaces $H$ and $H^{\prime}$ contained in $\partial\Omega$ with distinct linear parts in $\overline{\mathcal{F}}$. Then $H$ and $H^{\prime}$ are transverse, so they intersect at some point $y\in\partial\Omega$, which contradicts Lemma 3.5. Therefore $H$ and $H^{\prime}$ must have the same linear part; in other words $H$ and $H^{\prime}$ are parallel.
@@
Next, let $H_{0}$ denote the common linear part of the hyperplanes contained in $\partial\Omega$. Since $\Gamma$ preserves $\Omega$, the group $L(\Gamma)$ preserves $H_{0}$. Let $x\in\Omega$ and suppose that $(H_{0}+x)$ is not contained in $\Omega$. Then there exists $y\in(x+H_{0})\cap\partial\Omega$, and by the previous argument $(y+H_{0})\subset\partial\Omega$. Hence $x\in\partial\Omega$, a contradiction. This completes the proof. ∎

### v2

@@ 1.2.1. Reduction result
Fix once and for all a Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$. We say that $g\in\mathrm{SL}_{d}(\mathbb{R})$ does not contract a subspace $F\subset\mathbb{R}^{d}$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $x\in F$.
@@
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\mathsf{H}\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a group acting properly and cocompactly on $\Omega$. Let us denote by $L(\mathsf{H})$ the linear part of $\mathsf{H}$. Assume there exists $1\leq p\leq d-1$ and a closed $L(\mathsf{H})$-invariant subset $\mathcal{F}$ of $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, such that:
@@
    for all $g\in L(\mathsf{H})$, there exists an element of $\mathcal{F}$ which is not contracted by $g$;
@@
    If $x,y\in\mathcal{F}$ are distinct, then $x+y=\mathbb{R}^{d}$.
Then there exists $F$ in $\mathcal{F}$ that is $L(\mathsf{H})$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $F$.
In this proposition, the existence of an element in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ which is non-contracted by $L(\mathsf{H})$ suggests that $n-p\geq\operatorname{disc}(L(\mathsf{H}))$, where $\operatorname{disc}(L(\mathsf{H}))$ is the disompacity of $L(\mathsf{H})$ in the sense of Carrière [Car89]. In Carrière’s paper, the assumption of discompacity 1 is necessary. Proposition B says that, by adding a transversality assumption, we can relax this condition and get an intermediate structural result for groups of potentially higher dicsompacity. It generalizes a result of Tholozan [Tho15, Proposition 2.4], where the statement is proved in the case where $L(\mathsf{H})$ is contained in $\mathrm{U}(n,1)$ and $\mathcal{F}$ is the set of a complex hyperplanes.
In our case, the set of isotropic planes in $\mathbb{R}^{2,2}$ has two connected components, each of which is an orbit under the identity component ${\operatorname{SO}}_{0}(2,2)$; moreover, the elements of each orbit are pairwise transverse (see Proposition 3.9), meaning that if $x,y$ are two distinct isotropic planes in the orbit, then $x+y=\mathbb{R}^{4}$, i.e. they satisfy the second condition of Proposition B. Let us briefly explain this well known fact. Write $\mathbb{R}^{2,2}=\mathbb{R}^{2}\oplus\mathbb{R}^{2}$, endowed with the nondegenerate quadratic form of signature $(2,2)$ given by $g\oplus(-g)$, where $g$ is the standard Euclidean metric. The graph of any linear isometry $A\in O(2)$ is a totally isotropic plane in $\mathbb{R}^{2,2}$. One can show that the set of isotropic planes in $\mathbb{R}^{2,2}$ is exactly $O(2)$; in particular, it has two connected components. The key observation is that if $A,B\in{\operatorname{SO}}(2)$ are distinct elements, then their graphs intersect only at $0\in\mathbb{R}^{2,2}$ and so they are transverse. Indeed, their intersection coincides with the set of fixed vectors of $AB^{-1}$, and a nontrivial element of ${\operatorname{SO}}(2)$ has no nonzero fixed vectors. This property no longer holds in ${\operatorname{SO}}(d)$ for $d>2$, and it is precisely at this point that the signature $(2,2)$ plays a crucial role.
Nevertheless, Proposition B applies in other settings, including cases that deserve further investigation. For instance, when $d=8$ and $L(\mathsf{H})$ is contained in ${\operatorname{SO}}(4,\mathbb{C})$, or for any $d$ when $L(\mathsf{H})$ is contained in an irreducible simple rank-one Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$. A more general Lie-theoretic corollary of Proposition B, containing these two examples, is stated in Corollary 3.7. In another direction, an asymptotic version of Proposition B is part of a work in progress and applies to some cases where $L(\mathsf{H})$ is an $\{\alpha_{p}\}$-Anosov subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ in the sense of [Lab06, GW12].
@@ 1.2.2. Quotient geometry
Let $G$ be a homothety Lie group, and let $G_{\theta}=R\ltimes_{\theta}G$ be a semidirect product, where $R$ is a linear Lie group and $\theta$ commutes with a nontrivial homothety of $G$. Let $\pi\colon G_{\theta}\to R$ be the natural projection. Let $\Gamma\leq G_{\theta}$ be a discrete subgroup, and let $H=\overline{\pi(\Gamma)}$. Set $\Gamma_{nd}:=\Gamma\cap\pi^{-1}(H^{\circ})$. Then $\Gamma_{nd}$ admits a nilpotent syndetic hull $S$ in $G_{\theta}$ with the property that $\overline{\pi(\Gamma_{nd})}=\overline{\pi(S)}=H^{\circ}$.
@@ 2.2. Algebraic and cohomological constraints
Let $\Omega\subset\mathbb{R}^{d}$ be an open set divided by a discrete subgroup $\Gamma\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$. Then $\operatorname{cd}(\Gamma)\geq d$.
@@ 3.1. General reduction result
Fix once and for all the Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$. We say that an element $g$ of ${\operatorname{SL}}_{d}(\mathbb{R})$ does not contract $F$ with respect to $|\cdot|$ if $|g\cdot x|\geq|x|$ for all $g\in X$ and all $x\in F$. In what follows, we will simply say that $g$ does not contract $F$. To prove Proposition 3.1, we first establish the stronger result given by Proposition B, inspired by arguments from [Tho15]. To fix notation, let $p\in\llbracket 1,d\rrbracket$, and consider $\operatorname{Gr}_{p}(\mathbb{R}^{d})$, the Grassmannian of $p$-dimensional subspaces, note that the linear action of $\mathrm{SL}(d,\mathbb{R})$ induces a natural action on $\operatorname{Gr}_{p}(\mathbb{R}^{d})$.
The proof of Proposition B requires some preparation. Throughout this section, we work under the assumptions of Proposition B. Write $\Omega=\mathsf{H}\cdot\mathcal{C}$ with $\mathcal{C}\subset\Omega$ compact. Let $\varepsilon>0$ be such that $\mathcal{C}+B_{\varepsilon}\subset\Omega$, where $B_{\varepsilon}$ denotes the Euclidean ball of radius $\varepsilon$ centered at 0, defined with respect to the Euclidean norm $\lvert\cdot\rvert$.
@@
Let $y\in\partial\Omega$. Then there exist sequences $(g_{n})_{n\in\mathbb{N}}\subset\mathsf{H}$, $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$, and $(F_{n})_{n\in\mathbb{N}}\subset\mathcal{F}$ such that, up to extraction of a subsequence:
@@
    for every $n$, one has $(F_{n}+g_{n}\cdot x_{n})\cap B(g_{n}\cdot x,\varepsilon)\subset\Omega$;
@@
    $F_{n}$ converges in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ to some element $F^{y}$ in the closure of $\mathcal{F}$, with $(F^{y}+y)\cap B(y,\varepsilon)\subset\partial\Omega$, where $B(y,\varepsilon)$ is Euclidean ball of radius $\varepsilon$ centered at $y$.
@@
Since $\Omega=\mathsf{H}\cdot\mathcal{C}$, there exist sequences $(x_{n})_{n\in\mathbb{N}}\subset\mathcal{C}$ and $(g_{n})_{n\in\mathbb{N}}\subset\mathsf{H}$ such that $g_{n}\cdot x_{n}\to y$. This proves the first point. For all $n\in\mathbb{N}$, we have
@@
Let $\mathcal{F}$ be the subset of $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ appearing in the assumptions of Proposition B. For each $n$, there exists $F_{n}\in\mathcal{F}$ such that $g_{n}$ does not contract $F_{n}$, and set $F_{n}:=L(g_{n})\cdot F_{n}$. By the non-contracting assumption, we know that $L(g_{n})\cdot B_{\varepsilon}$ contains $F_{n}\cap B_{\varepsilon}$.
It follows from (1) that $\Omega$ contains $y_{n}+(F_{n}\cap B_{\varepsilon})$, establishing the second point. By compactness of $\mathcal{F}$, we may assume (up to extracting a subsequence) that $F_{n}$ converges in the Grassmannian to some $F^{y}\in\mathcal{F}$. Then $\overline{\Omega}$ contains $y+(F^{y}\cap B_{\varepsilon})$, and since $y\in\partial\Omega$, the inclusion must in fact be
$(F^{y}\cap B_{\varepsilon})+y\subset\partial\Omega.$
@@
In the previous lemma, we do not use the transversality of the elements of $\mathcal{F}$. In particular, the lemma asserts that each point of the boundary contains a small neighborhood of an affine subspace.
The next step is to show the uniqueness of the subspace $F^{y}$. Before doing so, we record the following elementary fact about transversality.
@@
Let $y\in\partial\Omega$, and let $F^{y}$ be an element of the closure of $\mathcal{F}$ as in Lemma 3.2. Then $F^{y}$ is the unique element of $\mathcal{F}$ with the property that there exists a neighborhood $\mathcal{V}$ of $y$ in $\Omega$ such that $(y+F^{y})\cap\mathcal{V}\subset\partial\Omega$.
@@
Let $(y_{n})_{n\in\mathbb{N}}$, $(F_{n})_{n\in\mathbb{N}}$, and $F^{y}$ be as in Lemma 3.2. By shrinking $\varepsilon$ if necessary, we may assume that $B(y,\varepsilon)\subset\mathcal{V}$.
We proceed by contradiction and assume the existence of a subspace $F\in\mathcal{F}$, distinct from $F^{y}$, that satisfies the hypothesis of the lemma. Then by assumption $(2)$ of Proposition B, one has $F^{y}+F=\mathbb{R}^{d}$.
Since $y_{n}\to y$ and $F_{n}\to F^{y}$, and using Lemma 3.4, we deduce that, for $n$ large enough, the set $(F_{n}+y_{n})\cap B(y_{n},\varepsilon)$ intersects $(F+y)\cap B(y,\varepsilon)$.
However, by Lemma 3.2, we have $(F_{n}+y_{n})\cap B(y_{n},\varepsilon)\subset\Omega$, while $(F+y)\cap B(y,\varepsilon)\subset(F+y)\cap\mathcal{V}\subset\partial\Omega$, a contradiction. ∎
@@
Let $y\in\partial\Omega$, and let $F^{y}$ be as in Proposition 3.2. Then $F^{y}+y\subset\partial\Omega$.
@@
Let $\mathcal{U}$ be the set of points $z\in F^{y}+y$ such that $(F^{y}+y)\cap B(z,\varepsilon)\subset\partial\Omega$. We will show that $\mathcal{U}$ is both open and closed (and nonempty) in $F^{y}+y$, so that $\mathcal{U}=F^{y}+y$, which will complete the proof.
Clearly, the set $\mathcal{U}$ is closed and contains $y$. We now show that $\mathcal{U}$ is also open. Let $z\in\mathcal{U}$ and $z^{\prime}\in(F^{y}+y)\cap B(z,\varepsilon/2)$. Then $(F^{y}+y)\cap B(z^{\prime},\varepsilon/2)\subset\partial\Omega$. Thus, by Lemma 3.5, we have $F^{y}=F^{z^{\prime}}$. By the definition of $F^{z^{\prime}}$, this implies $F^{y}\cap B(z^{\prime},\varepsilon)=F^{z^{\prime}}\cap B(z^{\prime},\varepsilon)\subset\partial\Omega,$ hence $z^{\prime}\in\mathcal{U}$. Therefore $\mathcal{U}$ is open, which completes the proof. ∎
@@
First, we claim that two affine subspaces contained in $\partial\Omega$ whose linear parts lie in $\mathcal{F}$ must be parallel. By Lemma 3.6, there exists at least one affine subspace contained in $\partial\Omega$ whose linear part belongs to $\mathcal{F}$.
@@
Assume, for contradiction, that there exist two affine subspaces $H$ and $H^{\prime}$ contained in $\partial\Omega$ with distinct linear parts in $\mathcal{F}$. Then $H$ and $H^{\prime}$ are transverse, so they intersect at some point $y\in\partial\Omega$, which contradicts Lemma 3.5. Therefore $H$ and $H^{\prime}$ must have the same linear part; in other words $H$ and $H^{\prime}$ are parallel.
@@
Next, let $H_{0}$ denote the common linear part of the hyperplanes contained in $\partial\Omega$. Since $\mathsf{H}$ preserves $\Omega$, the group $L(\mathsf{H})$ preserves $H_{0}$. Let $x\in\Omega$ and suppose that $(H_{0}+x)$ is not contained in $\Omega$. Then there exists $y\in(x+H_{0})\cap\partial\Omega$, and by the previous argument $(y+H_{0})\subset\partial\Omega$. Hence $x\in\partial\Omega$, a contradiction. This completes the proof. ∎
3.2. General reduction when the linear part is contained in a semisimple Lie subgroup
The goal of this section is to apply the general reduction result of Proposition B to a case of interest, namely when the linear part of $\mathsf{H}$ is contained in a semisimple Lie subgroup $G$ of ${\operatorname{SL}}_{d}(\mathbb{R})$; see Corollary 3.7 below. Indeed, in this case, the action of $G$ on the Grassmannians of $\mathbb{R}^{d}$ is well understood through the Cartan decomposition of $G$. This notion is recalled in next subsection .
Our goal in Section 3.3 will be to apply Corollary 3.7 to the case where $G={\operatorname{SO}}_{0}(2,2)$.
3.2.1. Cartan projection
Consider the subgroup ${\operatorname{SO}}(d)$ of ${\operatorname{SL}}_{d}(\mathbb{R})$ and the Cartan subspace of $\operatorname{Mat}(d,\mathbb{R})$ given by
$\overline{\mathfrak{a}}^{+}:=\{\operatorname{diag}(\lambda_{1},\dots,\lambda_{d})\mid\lambda_{1}\geq\dots\geq\lambda_{d}\ ,\ \ \sum_{k=1}^{d}\lambda_{k}=0\}.$
Then ${\operatorname{SL}}_{d}(\mathbb{R})={\operatorname{SO}}(d)\exp(\overline{\mathfrak{a}}^{+}){\operatorname{SO}}(d)$. This is called the Cartan decomposition of ${\operatorname{SL}}_{d}(\mathbb{R})$.
Given an element $g\in{\operatorname{SL}}_{d}(\mathbb{R})$, we can thus write $g=k\exp(X)k^{\prime}$ according to this decomposition. The elements $k,k^{\prime}\in K$ are not unique, but the element $X\in\overline{\mathfrak{a}}^{+}$ is. It is called the Cartan projection of $g$ and denoted by $\mu(g)$. We also consider the maps $\mu_{1},\dots,\mu_{d}:{\operatorname{SL}}_{d}(\mathbb{R})\rightarrow\mathbb{R}$, so that
$\mu(g)=\operatorname{diag}(\mu_{1}(g),\dots,\mu_{d}(g))\quad\forall g\in{\operatorname{SL}}_{d}(\mathbb{R}).$
3.2.2. A corollary of Proposition B
We can now prove the following Corollary 3.7 of Proposition B. For this, we take the notation of Subsection 3.2.1.
Corollary 3.7.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, different from $\mathbb{R}^{d}$, and let $\mathsf{H}\leq{\operatorname{SL}}_{d}(\mathbb{R})\ltimes\mathbb{R}^{d}$ be a group acting properly and cocompactly on $\Omega$. Let $G$ be a reductive Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ containing the linear part $L(\mathsf{H})$ of $\mathsf{H}$. Let
$p_{0}:=\operatorname{card}\{1\leq p\leq d\mid\mu_{i}(G)\subset\mathbb{R}_{+}\}.$
Assume that there exists $\frac{d}{2}\leq p\leq p_{0}$ the closure $\mathcal{F}$ of the $G$-orbit of $\operatorname{Span}(e_{1},\dots,e_{p})$ in $\operatorname{Gr}_{p}(\mathbb{R}^{d})$ consists of pairwise transverse subspaces, where $(e_{1},\dots,e_{d})$ is the canonical basis of $\mathbb{R}^{d}$.
Then there exists $F\in\mathcal{F}$ that is $L(\mathsf{H})$-invariant, and $\Omega$ is foliated by affine subspaces parallel to $F$.
Proof.
Since $G$ is a reductive subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$, there always exists a compact subgroup $K$ of ${\operatorname{SO}}(d)$ such that $G=K\exp(\mu(G))K$. Note that this decomposition does not always coincide with the Cartan decomposition of $G$. For all $g\in G$, we can write $g=waw^{\prime}$ according to the above Cartan decomposition and the assumption that $p\leq p_{0}$, and $g$ does not contract $(w^{\prime})^{-1}\operatorname{Span}(e_{1},\dots,e_{p})$. Then all the conditions of Proposition B are satisfied, and the corollary follows. ∎
Remark 3.8.
Corollary 3.7 gives an concrete examples where the conditions of Proposition B are satisfied, such as when $d=8$ and $L(\mathsf{H})$ is contained in ${\operatorname{SO}}(4,\mathbb{C})$, or for any $d$ if $L(\mathsf{H})$ is contained in an irreducible simple rank-one Lie subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$. However, it does not cover all the cases where Proposition B applies: an asymtotic version of it, which is part of work in progress of the three authors, applies if $L(\mathsf{H})$ is an $\{\alpha_{p}\}$-Anosov subgroup of ${\operatorname{SL}}_{d}(\mathbb{R})$ in the sense of [Lab06, GW12], and if its $\{\alpha_{p}\}$-limit set satisfies an asymptotic version of the non-contracting assumption of Proposition B. This contracting assumption is for instance satisfied whenever $L(\mathsf{H})$ is contained in ${\sf{Sp}}(d,\mathbb{R})$ for even $d\geq 4$ and $p=d/2$, or in ${\operatorname{SO}}(p,d-p)$ for $p\geq 2$ and $d-p\geq\max(p,3)$. Hence the geometric reduction still holds for these groups but does not follow from Corollary 3.7.